OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

OpenStudy (owlcoffee):

Apply distributive and you'll obtain: \[-3x-9=-3x-3+5\] and once you separate the variables from the constants: \[-3x+3x=9-3+5\]

OpenStudy (anonymous):

Lol im so bad at math, is the answer no solution because I got 3

OpenStudy (anonymous):

If its not 3 then i think its 6 lawl

OpenStudy (owlcoffee):

well if you operate you get \[0=11\] And since there is no "x" there, means that there is no solution to the equation.